OpenStudy (anonymous):

Solve, using the quadratic formula: 2x2 - 3x + 1 = 0 Answer {1, 0.5} {2.4, -0.9} {-1, -0.5} {-2.4, 0.9}

OpenStudy (anonymous):

Ok. now for the quadratic formula - comparing with ax^2+bx+c = 0, a = 2, b = -3 and c = 1 b2-4ac = (-3).(-3) -4.2.1 => 9-8 = 1 Hence root of b2-4ac = 1 Hence x = -b +/- 1/ 2a => 3 +/-1 / 4 => 4/4 or 2/4 => 1 or 0.5 is the answer.

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

yep no probs; just a new person here :)